# K3 surfaces associated to a cubic fourfold

### v1

@@ 1. Introduction
Let $\mathcal{C}$ be the moduli space of smooth projective cubic fourfolds and let $\mathcal{C}_{d}\subset\mathcal{C}$ be the Hasset divisor parametrizing special cubic fourfolds with a labelling of discriminant d, i.e. a positive defined rank-two primitive sublattice $K_{d}=<h^{2},v>\subset A(X)$. B. Hasset in [Hass ] proved that $X\in\mathcal{C}_{d}$ has an associated polarized K3 surface of degree $d$ if and only if $d$ satisfies the following numerical condition
@@
More precisely: for $d$ satisfying (*) $\mathcal{C}_{d}$ is birational to $\mathcal{N}_{d}$ if $d\equiv 2(6)$, where $\mathcal{N}_{d}$ is the moduli space of polarized K3 surfaces of degree d. Otherwise $\mathcal{C}_{d}$ is birational to a quotient of $\mathcal{N}_{d}$ by an involution, see [Hass Cor.25].
The results of several Authors (Hasset,Kuznetsov, Addington-Thomas) suggest that a cubic fourfold is rational if and only if it has an associated K3 surface. Kuznetsov conjectured that X has an associated K3 surface if and only if there exists a semiorthogonal decomposition of the derived category $D^{b}(X)$ of bounded complexes of coherent sheaves
@@
where $\mathcal{F}_{g}$ is the 19-dimensional moduli space of smooth polarized K3 surfaces $(S,H)$ of genus $g$, with $H^{2}=2g-2$and $g=n^{2}+n+2$.sending $(S,H)$ to the class $|X|$ of a cubic fourfold $X\in\mathcal{C}_{d}$. The surface $S$ is associated to the cubic fourfold $X$ and there is an isomorphism
@@
In [RS] it has been proved that a general $X\in\mathcal{C}_{d}$ with $d=14,26,38,42$ is rational. In all these cases the construction of a birational map $\mu:X\dashrightarrow W$, where $W$ is either $\mathbf{P}^{4}$ or a notable smooth rational fourfold, can be achieved by considering a surface $S_{d}\subset X$ admitting a four dimensional family of $(3e-1)$-secant curves of degree $e\geq 2$. These curves are parametrized by a rational variety with the property that trough a general point of $\mathbf{P}^{5}$ there passes a unique curve of the family, The inverse $\mu^{-1}:W\dashrightarrow X$ is given by a linear system of divisors in $|\mathcal{O}_{W}(e\cdot i(W)-1)|$, where $i(W)$ is the index of $W$, having points of multiplicity $e$ along an irreducible surface $U\subset W$ which is a birational incarnation of a K3 surface $S$ associated to $X$.
@@
where $\rho(S)$ is the rank of the Neron-Severi $N(SX$ ,see [KMP].
@@
The existence of an isomorphism as in 1.10 implies that the motive $h(X)$ is finite-dimensional, in the sense of [Kim], if and only if $h(S)$ is such. This has been proved for some families of cubic fourfolds, see [ABP] and [Lat]. In all these cases the motives $h(X)$ and $h(S)$ are of abelian type, i.e. they lie in subcategory $\mathcal{M}^{Ab}_{rat}(\mathbf{C})$ generated by the motives of abelian varieties.
@@
In this note we consider the cases of involutions and of prime order automorphisms. Note that a non-symplectic prime order automorphism can have order 2 or 3.
In Section 2 we show that for every cubic fourfold $X$ with an involution the motive $h(X)$ is associated to the motive of $S$ in the sense of 1.10. For two distinct families of cubic fourfolds $X$ with an involution, there is no K3 surface Hodge theoretically associated to $X$, while there is one family consisting of cubic fourfolds with an involution that have an associated K3 surface, are rational and their motives are of abelian type.
In Sect 3 we deal with cubic fourfolds $X$ admitting a non -symplectic automorphism of order 3 and prove that there cases where $X$ has an associated K3 surface and is rational as well as cases where $X$ is conjecturally irrational.
In Sect. 4 we describe families of special cubic fourfolds whose associated K3 surfaces are singular, i.e. their Picard rank is 20.
In Sect.5 we restrict to the case of cubic fourfolds $X$ such that $F(X)\simeq S^{2]}$, with $S$ a K3 surface, and show that the motive $h(X)$ is associated to the motive of $S$ in the sense of 1.10. We also show that a symplectic automorphism $\sigma$ on $S$ induces a symplectic automorphism $\tau$ on $X$ and give examples in the case $\sigma$ has order 3.
@@ 3. Automorphisms of order 3
with $F$ homogeneous of degree 3. The action of $\sigma_{1}$ is non -symplectic because $\det\sigma_{1}=\zeta$ and $\sigma(F)=F$.The cubic fourfold $X$ is a triple cover of $\mathbf{P}^{4}$ ramified along the cubic 3-fold $Y:(F(x_{0},x_{1},x_{2},x_{3},x_{4})=0)$.The dimension of the family $V_{1}$ is 10, see [GAL]. All these fourfolds have a motive of abelian type, see [Lat ].
@@
(1) the family $V_{1}$ contains a Pfaffian cubic fourfold $Y$ which contains no plane, see [BCS 2,Prop.5.1].Therefore $Y\in\mathcal{C}_{14},Y\notin\mathcal{C}_{8}$. There is an isomorphism $F(Y)\simeq S^{[2]}$, where $S$ is a K3 surface and an isomorphism $t_{2}(S)(1)\simeq t(Y)$. Since every cubic fourfold in $V_{1}$ has a motive of abelian type also $S$ has a motive of abelian type..
@@
which is isomorphic to $X$, see [CT, Prop.1.2].
@@
R.Laza in [Laz ] defined the algebraicity index $k_{X}$ of a smooth cubic fourfold $X$
$k_{X}=2^{\rho}/d_{X},$
where $\rho={\operatorname{rank}}A(X)_{prim}$, $d_{X}=det(A(X))$ and proved that $k_{X}\leq 1$ if $X$ has no Hodge theoretically associated K3 surface.
@@
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1 . Let $V_{3}$ be the family of cubic fourfolds with a non-symplectically action of the automorphism $\sigma_{3}$.The family $V_{3}$ has dimension 7. For a general element $X\in V_{3}$ the lattice $A(X)$ has rank 7 and $k_{X}<1$
@@
In [BCS 1, Ex.6.6] it is proved that the fixed sublattice $T_{\sigma_{3}}(X)$ for the action of automorphism induced by $\sigma_{3}$ on $F(X)$ is isomorphic to the lattice $<6>\oplus E^{*}_{6}(3)$. For $X$ general in $V_{3}$ the lattice $T_{\sigma_{3}}(X)$ coincides with $NS(F(X)$. From the isomorphism in 3.4 we get $A^{2}(X)_{prim}\simeq E^{*}_{6}(3)$. The lattice $L=E^{*}_{6}(3)$ is even, negative defined and 3-elementary with $A_{L}\simeq(\mathbf{Z}/3\mathbf{Z})^{5}$. Here $A_{L}=L^{*}/L$, where $L^{*}=\operatorname{Hom}(L,\mathbf{Z})$. Let $\{e_{i}\}_{1\leq i\leq 6}$ be a basis for $L$ and let $M=(<e_{i},e_{j}>)_{i,j}$ be the Gram matrix. Then $\det M$ equals the order of the group $A_{L}$ that is $2433^{5}=$ 243. Since $2^{\rho}=2^{6}$ we get $k_{X}<1$. ∎
@@
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1. There are two families of cubic fourfolds $X$ which are invariant under the action of $\sigma_{4}$, a first one where $\sigma_{4}$ acts non-symplectically and a second one where $\sigma_{4}$ acts symplectically, see [GAL,Thm. 3.8]. The family $V_{4}$ with a non-symplectical action of $\sigma$ has been described in [BCS 1, Ex.6.7]
Proposition 3.9.
The family $V_{4}$ has dimension 6. For a general element $X\in V_{4}$ the lattice $A(X)_{prim}$ has rank 9 and $X$ has no Hodge theoretically associated K3 surface
Proof.
Every $X\in V_{4}$ is defined by an equation of the form
@@
where $L_{2},M_{2},N_{2},P_{2}$ are homogeneous polynomial of degree 2 and $L_{1},M_{1},N_{1}$ are linear factors. The automorphism $\sigma_{4}$ acts non-symplectically on $F(X)$ because $\det\sigma_{4}=1$ and $\sigma^{*}_{4}(F)=\zeta$.The invariant sublattice $T_{\sigma_{4}}(F(X))\subset H^{2}(F,\mathbf{Z})$, for the action of $\sigma_{4}$ on the Fano variety $F(X)$, has rank 9 and equals $<-2>\oplus U(3)\oplus E^{*}_{6}(3)$, see [BCS 1 ,Ex. 6.7].
For a general element $X\in V_{4}$ the lattice $T_{\sigma_{4}}(F(X))$ equals $NS(F(X)$. From the isomorphism in 3.4 we get
$A^{2}(X)_{prim}\simeq U(3)\oplus E^{*}_{6}(3).$
The lattice $U(3)\oplus E^{*}_{6}(3)$ admits a primitive embedding in the K3 lattice $\Lambda_{K3}\simeq E^{2}_{8}\oplus U^{\oplus 3}$, see [AS,Table 2]. Therefore the transcendental lattice $T_{X}$ does not admit a primitive embedding in $\Lambda_{K3}$, and $X$ has no Hodge theoretically associated K3 surface, see [Laza , 1.7] . ∎
@@ 4. Singular K3 surfaces
The following result appears in [ABP ,Thm.1.1] : For $3\leq n\leq 20$,
@@
for $d_{k}\geq 8,d_{k}\equiv 0,2[6]$ and $d_{3},..,d_{n}=6\displaystyle{\prod_{i}}p_{i}^{2}$ or $26\displaystyle{\prod_{i}}p_{i}^{2}+$ 2 with $p_{i}$ a prime number.
@@
Let ${\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}\neq\emptyset$ with at least one of the divisor $\mathcal{C}_{d_{k}}$ parametrizing cubic fourfolds with an associate K3 surface. For every $X\in\displaystyle{\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}$ the lattice $A(X)$ of algebraic cycles has rank 21, hence the primitive cohomology of the associated K3 surface has rank 19 and the Neron-Severi group has rank 20.
@@
Therefore the Fano variety $F(X)$ is birational to the Hilbert scheme of length 2 subschemes of a Kummer surface.
A different way to construct cubic fourfolds with associated singular K3 surfaces is to consider a finite group $G$ of symplectic automorphisms acting on $X$ such that the coinvariant lattice $S_{G}(X)$ has maximal rank. The moduli space parametrizing cubic fourfolds with a symplectic group $G$ is $20-{\operatorname{rank}}S_{G}$. The maximal case, i.e. ${\operatorname{rank}}S_{G}=20$ correspond to cubic fourfolds $X$ such that the lattice of algebraic cycles $A(X)=<S_{G},h^{2}>$ , with $h$ the class of hyperplane section, has rank 21.All these cubics have an associated K3 surface $S$ with Picard rank equal to 20. Therefore all these cubic fourfolds have a motive of abelian type.
@@
and has a symplectic automorphism of order 7.It is i also invariant with respect to a anti-symplectic involution, see [LZ,Cor.6.9]. Therefore $X_{C}$ contains an Eckardt points and belongs to $\mathcal{C}_{8}\cap\mathcal{C}_{12}$, see 2.3.
@@ 5. K3 surfaces with a symplectic automorphism
The Klein cubic fourfold $X$, whose equation is of the form
@@
The non-symplectic automorphisms of order 3 on the cubic fourfolds belonging to the families $V_{1}$ and $V_{3}$ in Sect.3 are non-standard, see [BCS 1,Cor.7.6]
@@
An automorphism of $F(X)$ is polarized if and only if it is induced from an automorphism of the cubic fourfold $X$, see [LFu,Lemma1.2].
@@
Let $(S,L)$ be polarized K3 surface of genus $g$ and degree $d$, with $L^{2}=2g-2=d$ and $d=2(n^{2}+n+1)$. Let $X\in\mathcal{C}_{d}$ be the corresponding cubic fourfold in 5.1 . Let $\sigma$ be a symplectic automorphism of $S$ of prime order $p$ and let $\sigma^{[2]}$ be the induced automorphism on $S^{[2]}\simeq F(X)$. The natural automorphism $\sigma^{[2]}$ on $F(X)$ is polarized and hence it is induced by an automorphism $\tau$ on $X$. The fixed locus of the automorphism $\tau$ on $X$ consists of a finite number of isolated points.
@@
where $\mathbf{P}(T_{S})$ is the geometric projectivization of the tangent bundle $T_{S}$ For each $p\in S$ $\Delta_{p}$ is the rational curve consisting of 0-dimensional subschemes $\xi\in\Delta$ such that $\operatorname{supp}(\xi)=p$, see[FV,2.1].
@@
with $f=|L|\in H^{2}(S,\mathbf{Z})$, see [Hass,Thm. 29]. Since the symplectic automorphism of the polarized K3 surface $S$ fixes the classes $f$ and $\sigma^{[2]}$ fixes the class of $\Delta$ the induced automorphism fixes the polarization $g$. Therefore the natural automorphism $\sigma^{[2]}$ on $F(X)$ is polarized and hence it is induced by an automorphism $\tau$ of $X$.
@@
(2)$\xi_{p}$ with $\xi_{p}$ a non reduced subscheme of length 2 supported on $p$ with $\sigma(p)=p$, see [Bois 4.2 and Rk. 6].
The fixed locus of $\sigma$ on $S$ consists on a finite number of isolated points and and the fixed 1-dimensional subschemes $\xi_{p}$ in (2) correspond to $\Delta_{p}\subset S^{[2]}$ in 5.8with $\sigma(p)=p$.

### v2

@@ 1. Introduction
Let $\mathcal{C}$ be the moduli space of smooth projective cubic fourfolds and let $\mathcal{C}_{d}\subset\mathcal{C}$ be the Hassett divisor parametrizing special cubic fourfolds with a labelling of discriminant d, i.e. a positive defined rank-two primitive sublattice $K_{d}=<h^{2},v>\subset A(X)$. B. Hassett in [Hass] proved that $X\in\mathcal{C}_{d}$ has an associated polarized K3 surface of degree $d$ if and only if $d$ satisfies the following numerical condition
@@
More precisely: for $d$ satisfying (*) $\mathcal{C}_{d}$ is birational to $\mathcal{N}_{d}$ if $d\equiv 2(6)$, where $\mathcal{N}_{d}$ is the moduli space of polarized K3 surfaces of degree d. Otherwise $\mathcal{C}_{d}$ is birational to a quotient of $\mathcal{N}_{d}$ by an involution, see [Hass, Cor.25].
The results of several Authors (Hassett,Kuznetsov, Addington-Thomas) suggest that a cubic fourfold is rational if and only if it has an associated K3 surface. Kuznetsov conjectured that X has an associated K3 surface if and only if there exists a semi-orthogonal decomposition of the derived category $D^{b}(X)$ of bounded complexes of coherent sheaves
@@
where $\mathcal{F}_{g}$ is the 19-dimensional moduli space of smooth polarized K3 surfaces $(S,H)$ of genus $g$ with $H^{2}=2g-2$ and $g=n^{2}+n+2$. The map $\phi$ sends $(S,H)$ to the class $|X|$ of a cubic fourfold $X\in\mathcal{C}_{d}$. The surface $S$ is associated to the cubic fourfold $X$ and there is an isomorphism
@@
In [RS] it has been proved that a general $X\in\mathcal{C}_{d}$ with $d=14,26,38,42$ is rational. In all these cases the construction of a birational map $\mu:X\dashrightarrow W$, where $W$ is either $\mathbf{P}^{4}$ or a notable smooth rational fourfold, can be achieved by considering a surface $S_{d}\subset X$ admitting a four dimensional family of $(3e-1)$-secant curves of degree $e\geq 2$. These curves are parametrized by a rational variety with the property that trough a general point of $\mathbf{P}^{5}$ there passes a unique curve of the family. The inverse $\mu^{-1}:W\dashrightarrow X$ is given by a linear system of divisors in $|\mathcal{O}_{W}(e\cdot i(W)-1)|$, where $i(W)$ is the index of $W$, having points of multiplicity $e$ along an irreducible surface $U\subset W$, which is a birational incarnation of a K3 surface $S$ associated to $X$.
@@
where $\rho(S)$ is the rank of the Neron-Severi $N(SX)$, see [KMP].
@@
The existence of an isomorphism as in 1.10 implies that the motive $h(X)$ is finite-dimensional, in the sense of [Kim], if and only if $h(S)$ is such.
Some families of cubic fourfolds with a finite-dimensional motive have been described in [Lat 1] and [Lat 3]. In all these cases the motive $h(X)$ is of abelian type, i.e. lies in subcategory $\mathcal{M}^{Ab}_{rat}(\mathbf{C})$ generated by the motives of abelian varieties.
@@
In this note we consider families of cubic fourfolds admitting either an involution or a non-symplectic automorphism of prime order. Note that a non-symplectic prime order automorphism can have order 2 or 3.
In Section 2 we show that for every cubic fourfold $X$ with an involution the motive $h(X)$ is associated to the motive of a K3 surface $S$ in the sense of 1.10. For two distinct families of cubic fourfolds $X$ with an involution, there is no K3 surface Hodge theoretically associated to $X$, while there is one family consisting of cubic fourfolds with an involution that have an associated K3 surface, are rational and their motives are of abelian type.
In Sect 3 we deal with cubic fourfolds $X$ admitting a non -symplectic automorphism of order 3 and prove that there are 2 cases where $X$ has an associated K3 surface and is rational and 2 cases where $X$ is conjecturally irrational.
In Sect. 4 we describe families of special cubic fourfolds whose associated K3 surfaces are singular, i.e. their Picard rank is 20. These special cubic fourfolds have amotive of abelian type.
In Sect.5 we restrict to the case of cubic fourfolds $X$ such that $F(X)\simeq S^{2]}$, with $S$ a K3 surface, and show that the motive $h(X)$ is associated to the motive of $S$ in the sense of 1.10. We also show that a symplectic automorphism $\sigma$ on $S$ induces a symplectic automorphism $\tau$ on $X$ and give examples, in the case $\sigma$ has order 3, of K3 surfaces $S$ whose motive is associated to different cubic fourfolds $X_{1}\in\mathcal{C}_{d_{1}}$ and $X_{2}\in\mathcal{C}_{d_{2}}$ with $d_{1}\neq d_{2}$.
@@ 3. Automorphisms of order 3
with $F$ homogeneous of degree 3. The action of $\sigma_{1}$ is non -symplectic because $\det\sigma_{1}=\zeta$ and $\sigma(F)=F$.The cubic fourfold $X$ is a triple cover of $\mathbf{P}^{4}$ ramified along the cubic 3-fold $Y:(F(x_{0},x_{1},x_{2},x_{3},x_{4})=0)$.The dimension of the family $V_{1}$ is 10, see [GAL]. All these fourfolds have a motive of abelian type, see [Lat 1].
@@
(1) R.Laterveer in [Lat 2, Cor.5.2] proved that $A_{0}(Z)_{hom}=0$, with $Z=F(X)/<\sigma_{1}>$ and $X\in V_{1}$.
(2)Since ${\operatorname{rank}}A(X)=1$ the family $V_{1}$ contains contains no plane. The family $V_{1}$ contains a Pfaffian cubic fourfold $Y$ (see [BCS 2,Prop.5.1]. Since $Y\in\mathcal{C}_{14}$ there is an isomorphism $F(Y)\simeq S^{[2]}$, where $S$ is a K3 surface and an isomorphism $t_{2}(S)(1)\simeq t(Y)$. The K3 surface $S$ has a motive of abelian type.
@@
which is isomorphic to $X$, see [CT, Prop.1.2] and [Lat 3,Prop.6].
@@
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1 and let $V_{3}$ be the family of cubic fourfolds with a non-symplectically action of the automorphism $\sigma_{3}$.The family $V_{3}$ has dimension 7. For a general element $X\in V_{3}$ the lattice $A(X)$ has rank 7 and it is generated by the classes of 27 cubic scrolls. The fourfold $X$ belongs to the Hasset divisor $\mathcal{C}_{12}$, contains no plane and has no Hodge theoretically associated K3 surface. There is an isomorphism of motives in $\mathcal{M}_{rat}(\mathbf{C})$
$t_{2}(S)(1)\simeq t(X)$
where $S$ is a K3 surface.
@@
In [BG, Lemma 4.9 and Prop.4.11] it is proved that $X\notin\mathcal{C}_{d}$, for $d\equiv 2(6)$. Therefore $X$ contains no plane. The fourfold $X$ contains 27 families of cubic scrolls $\{[T_{i}],[T^{*}_{i}]\}$, $i=1,\cdots 27$, such that $A(X)$ is generated by $[T_{i}]$.Therefore $X\in\mathcal{C}_{12}$. Since $d=12$ satisfies the numerical condition in (**) there is an isomorphism of motives as in 1.9.
There is no primitive embedding of $T(X)(-1)$ in a K3 lattice, hence there is no associated K3 surface, see [BG,lemma 4.7].
Remark 3.9.
The family $V_{3}$ contains the cubic fourfold $Y$ defined by the equation
$Y:f(x_{0},x_{1},x_{2})+g(x_{3},x_{4})+x^{3}_{5}=0$
which is also invariant under the symplectic automorphism
$\tau:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},x_{2},\zeta x_{3},\zeta x_{4},\zeta x_{5}].$
The cubic fourfold $Y$ is rational and has a motive of abelian type, see [CT] and [Lat 3]. The fixed locus $\sigma_{3}$ on $F(X)$, for a general $X\in V_{3}$, is parametrized by $C\times D$, were $C,D$ are the elliptic curves defined by
$C:f(x_{0},x_{1},x_{2})=0\ in\ \mathbf{P}^{2}:x_{3}=x_{4}=x_{5}=0,$
$D:g(x_{3},x_{4})+x^{5}=0\ in\ \mathbf{P}^{2}:x_{0}=x_{1}=x_{2}=0$
and coincides with the fixed locus for the action of $\tau$ on $F(Y)$. In [Kaw,Prop.3.2] it is proved that the quotient variety $F(Y)/<\tau>$ has a crepant resolution $\tilde{Y}\to F(Y)/<\tau>$, with $\tilde{Y}$ birational to the generalized Kummer variety $K^{2}(C\times D)$. The same argument shows that the motive of the quotient $F(X)/<\sigma_{3}>$ is a direct summand of $K^{2}(C\times D)$, hence it is of abelian type.
@@
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1. There are two families of cubic fourfolds which are invariant under the action of $\sigma_{4}$, a first one where $\sigma_{4}$ acts non-symplectically and a second one where $\sigma_{4}$ acts symplectically, see [GAL,Thm. 3.8]. Let $V_{4}$ be the family of cubic fourfolds with a non-symplectical action of $\sigma_{4}$. Every $X\in X$ has an equation of the form
@@
where $L_{2},M_{2},N_{2},P_{2}$ are homogeneous polynomial of degree 2 and $L_{1},M_{1},N_{1}$ are linear factors. The automorphism $\sigma_{4}$ acts non-symplectically on $F(X)$ because $\det\sigma_{4}=1$ and $\sigma^{*}_{4}(F)=\zeta$.
Proposition 3.10.
The family $V_{4}$ has dimension 6. A general fourfold $X\in V_{4}$ contains nine disjoint planes $F_{1},\cdots,F_{9}$ and $A(X)$ has a basis given by $\{h^{2},[F_{1}],\cdots,[F_{8}]\}$. Therefore $X\in\mathcal{C}_{14}$, is rational and has an associated K3 surface.
Proof.
A plane in $\mathbf{P}^{5}$ that is invariant for the action of $\sigma_{4}$ has an equation of the form
$\{ax_{0}=bx_{1},cx_{2}=dx_{3},ex_{4}=fx_{5}\}.$
There are exactly nine disjoint planes $F_{1},\cdots,F_{9}$ that are invariant under $\sigma_{4}$ and are contained in $X$. Moreover
$h^{2}=(1/3)\sum_{1\leq i\leq 9}[F_{i}],$
see [BG, Prop. 4.13]. ∎
@@ 4. Singular K3 surfaces
The following result appears in [ABP ,Thm.1.1] : let $3\leq n\leq 20$ and let $d_{k}\geq 8,d_{k}\equiv 0,2[6]$, where
[…]
with $p_{i}$ a prime number. Then
@@
Let ${\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}\neq\emptyset$, with at least one of the divisor $\mathcal{C}_{d_{k}}$ parametrizing cubic fourfolds with an associate K3 surface. For every $X\in\displaystyle{\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}$ the lattice $A(X)$ of algebraic cycles has rank 21, hence the primitive cohomology of the associated K3 surface has rank 19 and the Neron-Severi group has rank 20.
@@
Therefore the motive $h(X)$ is associated to $h(S)$ and are both of abelian type. The Fano variety $F(X)$ is birational to the Hilbert scheme of length 2 subschemes of a Kummer surface.
A different way to construct cubic fourfolds with associated singular K3 surfaces is to consider a finite group $G$ of symplectic automorphisms acting on $X$ such that the covariant lattice $S_{G}(X)$ has maximal rank. The moduli space parametrizing cubic fourfolds with a symplectic group $G$ has dimension $20-{\operatorname{rank}}S_{G}$. The case ${\operatorname{rank}}S_{G}=20$ corresponds to cubic fourfolds $X$ such that the lattice of algebraic cycles $A(X)=<S_{G},h^{2}>$ , with $h$ the class of hyperplane section, has rank 21. All these cubics have an associated K3 surface $S$ with Picard rank equal to 20. Therefore all these cubic fourfolds have a motive of abelian type.
@@
and has a symplectic automorphism of order 7.It is i also invariant with respect to a anti-symplectic involution, see [LZ,Cor.6.9]. Therefore $X_{C}$ contains an Eckardt points and belongs to $\mathcal{C}_{8}\cap\mathcal{C}_{12}$.
@@ 5. K3 surfaces with a symplectic automorphism
Example 5.6.
(1) The Klein cubic fourfold $X$, whose equation is of the form
@@
(2) The non-symplectic automorphisms of order 3 on the cubic fourfolds belonging to the families $V_{1}$ and $V_{3}$ in Sect.3 are non-standard, see [BCS 1,Cor.7.6]
@@
An automorphism of $F(X)$ is polarized if and only if it is induced from an automorphism of the cubic fourfold $X$, see [Fu,Lemma1.2].
@@
Let $(S,L)$ be a polarized K3 surface of genus $g$ and degree $d$, with $L^{2}=2g-2=d$ and $d=2(n^{2}+n+1)$. Let $X\in\mathcal{C}_{d}$ be the corresponding cubic fourfold in 5.1. Let $\sigma$ be a symplectic automorphism of $S$ of prime order $p$ and let $\sigma^{[2]}$ be the induced automorphism on $S^{[2]}\simeq F(X)$. The natural automorphism $\sigma^{[2]}$ on $F(X)$ is polarized and hence it is induced by an automorphism $\tau$ on $X$. The fixed locus of the automorphism $\tau$ on $X$ consists of a finite number of isolated points.
@@
where $\mathbf{P}(T_{S})$ is the geometric projectivization of the tangent bundle $T_{S}$. For each $p\in S$ $\Delta_{p}$ is the rational curve consisting of 0-dimensional subschemes $\xi\in\Delta$ such that $\operatorname{supp}(\xi)=p$, see [FV,2.1].
@@
with $f=|L|\in H^{2}(S,\mathbf{Z})$, see [Hass, Thm.29]. The symplectic automorphism $\sigma$ of the polarized K3 surface $S$ fixes the classes $f$ and the automorphism $\sigma^{[2]}$ of $S^{[2]}$ fixes the class of $\Delta$. Therefore the induced automorphism fixes the polarization $g$. Therefore the natural automorphism $\sigma^{[2]}$ on $F(X)$ is polarized and hence it is induced by an automorphism $\tau$ of $X$.
@@
(2)$\xi_{p}$ with $\xi_{p}$ a non reduced subscheme of length 2 supported on $p$ with $\sigma(p)=p$, see [Bois, 4.2 and Rk. 6].
The fixed locus of $\sigma$ on $S$ consists on a finite number of isolated points and and the fixed 1-dimensional subschemes $\xi_{p}$ in (2) correspond to $\Delta_{p}\subset S^{[2]}$, with $\sigma(p)=p$.